# vCache: Verified Semantic Prompt Caching

### v3

@@ C.3 Threshold Dilemma
In the top plot of Figure 2, we separate all observations into two sets: one where $c(x)=1$ (correct cache hit) and another where $c(x)=0$ (incorrect hit). We plot the kernel density functions (KDF) to visualize the distribution of correct and incorrect observations. The result shows that correct and incorrect observations are nearly indistinguishable in similarity space, with substantial overlap and similar means (0.84 vs. 0.85). This illustrates that one single similarity threshold is not a reliable decision boundary.
@@
Appendix D Benchmark Creation
@@
D.1 SemCacheClassification Benchmark
The SemCacheClassification benchmark is designed to evaluate semantic caching in structured classification settings, such as those found in modern database environments [Liu et al., 2024b]. Several database systems—including Databricks, Snowflake, and AWS—have introduced LLM-based extensions to SQL via User-Defined Functions (UDFs), enabling capabilities beyond traditional SQL semantics [Liu et al., 2024b]. However, LLM-based UDFs are inherently non-deterministic, execute row-by-row, and pose integration challenges for parallelization and query optimization [Franz et al., 2024]. When table rows contain semantically similar values, semantic caching can reduce LLM invocation frequency, improving both latency and cost. This benchmark captures such use cases, where small variations in phrasing should not require repeated inference.
The benchmark consists of 45,000 short-form prompts with a fixed output label space. Each example follows a prompt–response format, where the prompt expresses a classification query and the expected response is a one-word label. The benchmark combines three diverse classification datasets: CommonsenseQA [Talmor et al., 2018], Ecommerce Categorization [Saurabh Shahane, 2023], and Amazon Instant Video Review [Ni et al., 2019]. Sample prompts and response formats from each dataset are shown below, and Table 1 summarizes label distributions across the benchmark.
A sample entry from the Ecommerce Categorization [Saurabh Shahane, 2023] dataset:
@@ D.1 SemCacheClassification Benchmark
A sample entry from the Commonsense QA Talmor et al. [2018] dataset:
@@
A sample entry from the Amazon Instant Video Review Ni et al. [2019] dataset:
@@
D.2 SemCacheLMArena Benchmark
The SemCacheLMArena benchmark is designed to evaluate semantic caching in chatbot environments, where users may issue semantically similar prompts with different phrasing. In such settings, caches must generalize across diverse surface forms while maintaining response correctness. This benchmark captures these challenges by grouping semantically similar user inputs and testing whether caching systems can reuse responses accurately.
To construct the benchmark, we use the LM-Arena human preference dataset [Zheng et al., 2023], which contains 100,000 real-world user queries. We randomly sample 3,500 distinct prompts, each of which defines a class. For each class, we generate between 1 and 23 semantically similar variants using GPT-4.1-nano, resulting in a total of 60,000 prompts. A class ID is assigned to each prompt to evaluate caching correctness: a cache hit is considered correct if the retrieved response belongs to the same class as the query. Figure 9 shows the distribution of class sizes (number of prompts belonging to a class), confirming broad variability in prompt paraphrasing. To support model-agnostic evaluation, we generate responses for all prompts using GPT-4.1-nano and GPT-4o-mini. The corresponding response length distributions are shown in Figure 10.
@@ D.2 SemCacheLMArena Benchmark
D.3 SemCacheSearchQueries Benchmark
The SemCacheSearchQueries benchmark is designed to evaluate semantic caching in open-domain search applications. Large-scale search engines, such as Google, increasingly rely on LLMs to generate direct answers to natural language queries [Liu et al., 2024a, Wang et al., 2024]. While this improves user experience, it introduces significant latency and cost, particularly at the scale of millions of daily queries. Many queries issued to search engines are paraphrased variations of earlier inputs, making semantic caching a natural fit for reducing redundant LLM inference in this setting.
The benchmark is constructed from a filtered subset of the ORCAS dataset [Craswell et al., 2020], containing real-world search engine queries. We begin by sampling 500,000 queries and embedding each using the gte-large-en-v1.5 embedding model [Zhang et al., 2024]. We then apply $k$-means clustering to group similar queries and retain the largest clusters, resulting in 150,000 entries. Within each cluster, we apply a union-find algorithm guided by an LLM-based judge (GPT-4.1-nano) to determine whether query pairs yield the same response. Sub-clusters identified in this step define the equivalence classes used for caching evaluation. Figure 11 summarizes the benchmark properties, including class size distribution, frequent query terms, and statistics on the number of queries per class.

### v4

@@ D.6 Threshold Dilemma
In the top plot of Figure 3, we separate all observations into two sets: one where $c(x)=1$ (correct cache hit) and another where $c(x)=0$ (incorrect hit). We plot the kernel density functions (KDF), also known as kernel density estimations (KDE), to visualize the distribution of correct and incorrect observations. The result shows that correct and incorrect observations are nearly indistinguishable in similarity space, with substantial overlap and similar means (0.84 vs. 0.85). This illustrates that one single similarity threshold is not a reliable decision boundary.
@@
Appendix E Benchmark Creation
@@
E.1 SemCacheClassification Benchmark
The SemCacheClassification benchmark is designed to evaluate semantic caching in structured classification settings, such as those found in modern database environments (Liu et al., 2024b). Several database systems, including Databricks, Snowflake, and AWS, have introduced LLM-based extensions to SQL via User-Defined Functions (UDFs), enabling capabilities beyond traditional SQL semantics (Liu et al., 2024b). However, LLM-based UDFs are inherently non-deterministic, execute row by row, and pose integration challenges for parallelization and query optimization (Franz et al., 2024). When table rows contain semantically similar values, semantic caching can reduce the frequency of LLM invocations, thereby improving both latency and cost. This benchmark captures such use cases, where slight variations in phrasing should not require repeated inference.
The benchmark consists of 45,000 short-form prompts with a fixed output label space. Each example follows a prompt–response format, where the prompt expresses a classification query and the expected response is a one-word label. The benchmark combines three diverse classification datasets: CommonsenseQA (Talmor et al., 2018), Ecommerce Categorization (Saurabh Shahane, 2023), and Amazon Instant Video Review (Ni et al., 2019). This dataset composition models out-of-distribution data because the three sources differ significantly in domain, style, and vocabulary, forcing semantic caching methods to generalize beyond a single homogeneous dataset. Sample prompts and response formats from each dataset are shown below, and Table 4 summarizes label distributions across the benchmark.
A sample entry from the Ecommerce Categorization (Saurabh Shahane, 2023) dataset:
@@ E.1 SemCacheClassification Benchmark
A sample entry from the Commonsense QA Talmor et al. (2018) dataset:
@@
A sample entry from the Amazon Instant Video Review Ni et al. (2019) dataset:
@@
E.2 SemCacheLMArena Benchmark
The SemCacheLMArena benchmark is designed to evaluate semantic caching in chatbot environments, where users may issue semantically similar prompts with different phrasing. In such settings, caches must generalize across diverse surface forms while maintaining response correctness. This benchmark addresses these challenges by grouping semantically similar user inputs and testing whether caching systems can accurately reuse responses.
To construct the benchmark, we use the LM-Arena human preference dataset (Zheng et al., 2023a), which contains 100,000 real-world user queries. We randomly sample 3,500 distinct prompts, each of which defines a class. For each class, we generate between 1 and 23 semantically similar variants using GPT-4.1-nano, resulting in a total of 60,000 prompts. A class ID is assigned to each prompt to evaluate caching correctness: a cache hit is considered correct if the retrieved response belongs to the same class as the query. Figure 10 shows the distribution of class sizes (number of prompts belonging to a class), confirming broad variability in prompt paraphrasing. To support model-agnostic evaluation, we generate responses for all prompts using GPT-4.1-nano and GPT-4o-mini. The corresponding response length distributions are shown in Figure 11.
@@ E.2 SemCacheLMArena Benchmark
E.3 SemCacheSearchQueries Benchmark
The SemCacheSearchQueries benchmark is designed to evaluate semantic caching in open-domain search applications. Large-scale search engines, such as Google, increasingly rely on LLMs to generate direct answers to natural language queries (Liu et al., 2024a; Wang et al., 2024). While this improves user experience, it introduces significant latency and cost, particularly at the scale of millions of daily queries. Many queries issued to search engines are paraphrased variations of earlier inputs, making semantic caching a natural fit for reducing redundant LLM inference in this setting.
The benchmark is constructed from a filtered subset of the ORCAS dataset (Craswell et al., 2020), containing real-world search engine queries. We begin by sampling 500,000 queries and embedding each using the gte-large-en-v1.5 embedding model (Zhang et al., 2024). We then apply $k$-means clustering to group similar queries and retain the largest clusters, resulting in 150,000 entries. Within each cluster, we apply a union-find algorithm guided by an LLM-based judge (GPT-4.1-nano) to determine whether query pairs yield the same response. Sub-clusters identified in this step define the equivalence classes used for caching evaluation. Figure 12 summarizes the benchmark properties, including class size distribution, frequent query terms, and statistics on the number of queries per class.
@@ E.3 SemCacheSearchQueries Benchmark
Appendix F Hyperparameter Guidance
vCache is designed to be simple to configure, with the error rate bound $\delta$ as its primary hyperparameter. We recommend setting this bound based on the desired trade-off between accuracy and cost. For high-accuracy applications (e.g., customer support or safety-critical systems), a conservative value such as 0.5% can be appropriate. For use cases with higher tolerance for occasional errors and stronger cost or latency constraints (e.g., search or summarization), values around 2–3% may be reasonable. Ultimately, the choice depends on application-specific requirements.